OpenStudy (anonymous):

-5/6 divided by 1/4= -3 1/3 but can it be simplified? explain

OpenStudy (anonymous):

no its should just be -3/1/3

OpenStudy (anonymous):

so if i put it then it will be right?

OpenStudy (anonymous):

it says to simplify though and that doesnt look right hmm

TheSmartOne (thesmartone):

\[-\frac{ 5 }{6 }\div \frac{ 1 }{4 }=\frac{ -5 }{ 6}\times \frac{ 4 }{ 1 }=\frac{ -5\times4 }{ 6\times1 }=\frac{ -20 }{6 }=\frac{ -10 }{ 3 }=-3\frac{ 1 }{ 3 }\]

OpenStudy (anonymous):

so i was right?

TheSmartOne (thesmartone):

yes...

OpenStudy (anonymous):

yay :3

OpenStudy (anonymous):

hold on

OpenStudy (anonymous):

lol nvm it will be a while before i need any help im watching a tutorial for this thing